assuming that Huawei produces 5% of their total production of smartphones are defective. If 10 items are to be chosen at random from the production line, what then is the probability that;


a. all smartphones are not defective?


b. at most 2 of the smartphones are defective?


c. at least 3 of the smartphones are defective?

1
Expert's answer
2022-03-01T08:19:29-0500

Let "X=" the number of defective smartphones: "X\sim Bin(n, p)."

Given "n=10, p=0.05, q=0.95."

a.


"P(X=0)=\dbinom{10}{0}(0.05)^0(0.95)^{10-0}"

"=0.59873693924"

b.


"P(X\le 2)=P(X=0)+P(X=1)+P(X=2)"

"=\dbinom{10}{0}(0.05)^0(0.95)^{10-0}"

"+\dbinom{10}{1}(0.05)^1(0.95)^{10-1}+\dbinom{10}{2}(0.05)^2(0.95)^{10-2}"

"=0.59873693924+0.31512470486"

"+0.07463479852=0.98849644262"

c.


"P(X\ge 3)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dbinom{10}{0}(0.05)^0(0.95)^{10-0}"

"-\dbinom{10}{1}(0.05)^1(0.95)^{10-1}-\dbinom{10}{2}(0.05)^2(0.95)^{10-2}"

"=1-0.59873693924-0.31512470486"

"-0.07463479852=0.01150355738"